OpenStudy (anonymous):

What is the greatest common factor of 16,32,and 44

OpenStudy (anonymous):

Do you know what the greatest common factor of a set of numbers is defined as?

OpenStudy (anonymous):

This sort of question can be brute-forced easily by simply writing out all of the factors of the numbers in question, at least for simple cases like this.

OpenStudy (solomonzelman):

\[16=4 \times\color{blue}{ 4 }\]\[32=8 \times \color{blue}{ 4 }\]\[44=11 \times \color{blue}{ 4 }\]

OpenStudy (solomonzelman):

Which factor they share in common?

OpenStudy (anonymous):

(couldn't possibly be the one you highlighted in blue)

OpenStudy (anonymous):

so it would be the highest number that goes into all the numbers?

OpenStudy (solomonzelman):

YES!

OpenStudy (anonymous):

thanks